### v1

# Modular invariance groups and defect McKay-Thompson series

Abstract
It has been known since 1992 that the McKay-Thompson series $T_{g}(q)$ of the Moonshine module form Hauptmoduln for genus zero subgroups of $SL(2,\mathbb{R})$ [1, 2, 3, 4]. In [5], Lin and Shao constructed a series analogous to the McKay-Thompson series (a twined partition function of the Monster CFT), but using a non-invertible topological defect rather than an element of the Monster group $\mathcal{M}$. This “defect McKay-Thompson series” was found to be invariant under a genus zero subgroup of $SL(2,\mathbb{R})$, but was shown not to be the Hauptmodul of the subgroup. Nevertheless, one might wonder if a weaker version of Borcherds’ theorem holds for non-invertible defects: perhaps defect McKay-Thompson series enjoy genus zero invariance groups in $SL(2,\mathbb{R})$, whether or not they are Hauptmoduln for those groups. Using the decompositions of the monster stress tensor found in [6], we construct several new defect McKay-Thompson series, study their modular properties, and determine their invariance groups in $SL(2,\mathbb{R})$. We discover that many of the invariance groups are not genus zero.
1 Introduction
The phenomenon of Monstrous Moonshine, conjectured by Conway and Norton in [1] and eventually proved by Borcherds in [4], connects the representation theory of the Monster group $\mathcal{M}$ with the theory of modular functions. Specifically, Conway and Norton conjectured that the McKay-Thompson series $T_{g}(q)$ associated with elements $g$ of $\mathcal{M}$ are Hauptmoduln for genus zero subgroups of $SL(2,\mathbb{R})$. Frenkel, Lepowsky, and Meurman’s construction in [3] of a vertex operator algebra (VOA) with outer automorphism group $\mathcal{M}$ established a connection between moonshine and conformal field theory. This remarkable correspondence has since led to deep insights into both number theory and the theory of VOAs.
In [5], Lin and Shao extended these ideas by introducing the “defect McKay-Thompson series.” Instead of computing the partition function of the monster CFT twined by a group element $g$, they considered the insertion of (generically noninvertible) topological defect lines (TDLs) into the partition function. Lin and Shao found an example of a noninvertible TDL in the monster CFT (the Kramers-Wannier defect), and showed that the associated defect McKay-Thompson series enjoyed a genus zero invariance group in $SL(2,\mathbb{R})$, though it was not the Hauptmodul. This work left open the question of whether all defect McKay-Thompson series share the genus-zero property of their classical counterparts or whether the genus-zero property found in [5] was a coincidence due to the modular functions being invariant under subgroups of low level.
In this work, we construct a number of new defect McKay-Thompson series, using decompositions of the stress tensor of the monster CFT into those of smaller RCFTs. These decompositions, inspired by Norton’s monstralizer pairs, the work of Höhn [7] and subsequent generalizations which are explicated and generalized in [6], allow us to (non-exhaustively) catalog TDLs of the monster theory, as well as their eigenvalues, from which we may construct the defect McKay-Thompson series. We may then use the known modular properties of the smaller RCFTs appearing in our decompositions to infer the modular properties of the defect McKay-Thompson series. Doing so, we find that the genus-zero property of the classical McKay-Thompson series does not in general seem to hold for their noninvertible counterparts.
The paper is organized as follows: In Section 2, we describe the methodology used to construct and analyze the defect McKay-Thompson series, including the use of monstralizer pairs. Section 3 describes our method for computing the modular invariance groups of the resulting objects, first in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ and then considering enlargements in $\operatorname{\textsl{SL}}(2,\mathbb{R})$. Finally, section 4 presents our findings on the invariance groups of several defect McKay-Thompson series, detailing both the genus zero and non-genus zero cases, and discusses potential extensions to this work.
2 TDLs from monstralizers
The term “monstralizer pair” is due to Norton [8], who used it to describe subgroups of $\mathcal{M}$ which mutually centralize one another. In [6], Bae et al. used Norton’s monstralizer pairs as a clue to construct decompositions of the stress tensor of the monster CFT in terms of smaller RCFTs whose symmetry groups are those subgroups of $\mathcal{M}$ appearing in the monstralizer pair. This definition may be thought of as the “VOA uplift” of the group theoretic definition of monstralizers. More precisely, if $T(z)$ is the stress tensor of the monster CFT, and CFTs $\mathcal{A}$ and $\mathcal{B}$ have symmetry groups which mutually centralize one another in $\mathcal{M}$, and stress tensors $t^{(\mathcal{A})}(z)$ and $t^{(\mathcal{B})}(z)$ satisfying
$T(z)=t^{(\mathcal{A})}(z)+t^{(\mathcal{B})}(z),$ (2.1)
then $\mathcal{A}$ and $\mathcal{B}$ may be said to form a “monstralizer pair at the level of CFTs,” and the Virasoro primaries of $V^{\natural}$ may be organized into sets of primaries with respect to $t^{(\mathcal{A})}$ or $t^{(\mathcal{B})}$ individually. This in turn implies that the partition function of $V^{\natural}$ may be decomposed as
$J(\tau)=\sum_{i}\chi_{i}^{(\mathcal{A})}(\tau)\chi_{i}^{(\mathcal{B})}(\tau)$ (2.2)
for some (possibly not all) characters $\chi_{i}^{(\mathcal{A})}$ of $A$ and $\chi_{i}^{(\mathcal{B})}$ of $B$. The modular $S$ and $T$ matrices of $\mathcal{A}$ and $\mathcal{B}$ are related by (2.2): precisely, we must have
$S^{(\mathcal{B})}=S^{(\mathcal{A})*}\mbox{\quad and\quad}T^{(\mathcal{B})}=T^{(\mathcal{A})*}$ (2.3)
in order that (2.2) be modular-invariant. A variety of monstralizer pairs of VOAs are cataloged in [6]. In each of those examples one of the CFTs is an RCFT, in fact a finite sum of minimal models.
Given a monstralizer pair, suppose that the CFT $\mathcal{A}$ contains a topological defect line $\mathcal{L}^{(\mathcal{A})}$, whose action on the Virasoro primaries of $\mathcal{A}$ is
$\mathcal{L}^{(\mathcal{A})}\phi_{i}^{(A)}=L_{i}^{(\mathcal{A})}\phi_{i}^{(\mathcal{A})}$ (2.4)
(all the TDLs that we will consider in this work act diagonally on the primaries of interest). Then, using (2.1) and the map between Virasoro primaries of $V^{\natural}$ and Virasoro primaries of $\mathcal{A}$ (call it $\pi$), we may construct a TDL $\mathcal{L}$ of the monster CFT by imposing that the action of $\mathcal{L}$ on a given Virasoro primary of $V^{\natural}$ (say $\phi_{i}$) is
$\mathcal{L}\phi_{i}=L_{\pi(\phi_{i})}^{(\mathcal{A})}\phi_{i}.$ (2.5)
Such a line commutes with $t^{(\mathcal{B})}$, since its action on the primaries of $t^{(\mathcal{B})}$ is diagonal, and commutes with $t^{(\mathcal{A})}$ since it acts on the primaries of $t^{(\mathcal{A})}$ identically to $\mathcal{L}^{(\mathcal{A})}$, which was by supposition a topological defect of $\mathcal{A}$. Thus, the line commutes with $T(z)$, so is topological, and forms a topological defect of $V^{\natural}$. We emphasize, though, that while $\mathcal{L}$ and $\mathcal{L}^{(\mathcal{A})}$ have the same eigenvalues, they may differ in other respects – for example, $\mathcal{L}$ is generically not a Verlinde line, whether or not $\mathcal{L}^{(\mathcal{A})}$ is. At the level of characters, this construction allows us to compute the partition function of the monster CFT twined by the line $\mathcal{L}$:
$Z^{\mathcal{L}}(\tau)=\sum_{i}L_{i}^{(\mathcal{A})}\chi_{i}^{(\mathcal{A})}(\tau)\chi_{i}^{(\mathcal{B})}(\tau)=\vec{\chi}^{(\mathcal{A})}(\tau)^{\intercal}D_{\mathcal{L}}\vec{\chi}^{(\mathcal{B})}(\tau)$ (2.6)
where $\vec{\chi}^{(\mathcal{A})}(\tau)$ is a vector formed of those characters of $\mathcal{A}$ which appear in the monstralizer decomposition (2.2), and $D_{\mathcal{L}}=\operatorname{\text{diag}}(L_{i}^{(\mathcal{A})})$. The objects $Z^{\mathcal{L}}(\tau)$ are known as “defect McKay-Thompson series,” a term defined in [5]. In that paper, Lin and Shao identified two TDLs of the monster theory, one corresponding to the $\mathbb{Z}_{2\text{A}}$ conjugacy class of the monster and one non-invertible line, analogous to the Kramers-Wannier defect of the Ising CFT. Using the monstralizer pair $\quantity(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$ in the notation of [6], we recover the same lines and the associated defect McKay-Thompson series (given in detail in section 2.1). We also consider the monstralizer pairs $\quantity(\mathbb{Z}_{3\text{A}},3.Fi^{\prime}_{24})$, $\quantity(\mathbb{Z}_{4\text{A}},4.2^{22}.Co_{3})$, and $\quantity(D_{3\text{A}},Fi_{23})$. In each case we discover new TDLs of the monster theory, and in all but one we are able to show that we have found all TDLs arising from the monstralizer pair under consideration.
2.1 $\quantity(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$: Ising CFT
The monstralizer pair $(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$ has been known at the level of CFTs since [7], though in different language. In that work, Höhn constructed a subVOA of $V^{\natural}$ whose automorphism group is the baby monster $\mathbb{B}$. Lin and Shao provided a physical explanation in [5]: the fermionization of the monster CFT under an element of the $\mathbb{Z}_{2\text{A}}$ conjugacy class is the direct product of Höhn’s baby monster VOA with a single Majorana-Weyl fermion. Conversely, under bosonization, they showed that the $\mathbb{Z}_{2}$ symmetry of the direct product theory corresponding to fermion parity becomes the non-invertible element of a fusion subcategory of the monster which is isomorphic to the Ising category. In our language, this subcategory arises from the monstralizer pair $\quantity(\mathbb{Z}_{2\text{A}},2.\mathbb{B})$, which uplift to VOAs as the Ising CFT and the baby monster CFT. The Verlinde lines of the Ising CFT exhaust the fusion category of TDLs, and are reproduced (along with their action on the primaries) for convenience in table 1. Note that in this table, as well as those to follow, the top row does not label individual primaries of the monster; rather it should be understood to label sets of Virasoro primaries of $V^{\natural}$ with respect to the Ising stress tensor contained within the monster stress tensor (cf. eq 2.1).
$1$ $\sigma$ $\epsilon$
$\eta$ $1$ $1$ $-1$
$\mathcal{D}$ $\sqrt{2}$ $-\sqrt{2}$ 00
Table 1: TDLs of the Ising subcategory within the monster CFT. We emphasize that the labels $1$, $\sigma$, $\epsilon$, which refer to individual primaries in the constext of the Ising CFT, here refer to sets of Virasoro primaries of the moonshine module which are also Virasoro primaries under the Ising factor in the decomposition of the monster stress tensor.
The $J$-invariant may be written in terms of Ising and baby monster characters as
$J(\tau)=\chi_{1}(\tau)\chi^{(V\mathbb{B}^{\natural})}_{1}(\tau)+\chi_{\sigma}(\tau)\chi^{(V\mathbb{B}^{\natural})}_{\sigma}(\tau)+\chi_{\epsilon}(\tau)\chi^{(V\mathbb{B}^{\natural})}_{\epsilon}(\tau)$ (2.7)
where $\quantity{1,\sigma,\epsilon}$ refer to the Ising primaries with $h=\quantity{0,\frac{1}{16},\frac{1}{2}}$ respectively, and the baby monster characters $\chi^{(V\mathbb{B}^{\natural})}_{i}$ are related to the Ising characters by the Hecke operator for vector-valued modular forms $\mathsf{T}_{47}$ [6, 9]. To obtain the defect McKay-Thompson series, we make use of equation 2.6; we find $q$-expansions for the series $Z^{\eta}$ and $Z^{\mathcal{D}}$ which agree with those in [5]. The line $\eta$ simply implements the $\mathbb{Z}_{2\text{A}}$ global symmetry of the monster, so $Z^{\eta}$ is a “typical” McKay-Thompson series, of the type studied by Borcherds, and satisfies the genus zero property. On the other hand, the line $\mathcal{D}$ is non-invertible, so $Z^{\mathcal{D}}$ is not of this type, and is not guaranteed to be genus zero.
2.2 $\quantity(\mathbb{Z}_{3\text{A}},3.Fi^{\prime}_{24})$: Three-state Potts model
The next monstralizer pair we consider is $(\mathbb{Z}_{3\text{A}},3.Fi^{\prime}_{24})$, which uplift to VOAs as the three-state Potts model and its commutant in $V^{\natural}$. Unlike the Ising case, the commutant VOA here is not a familiar one (though it was studied in [10]); we shall only need its characters, which are images of the three-state Potts characters under the action of the Hecke operator $\mathsf{T}_{29}$. Also unlike the Ising case, the Verlinde lines of the three-state Potts model do not exhaust its fusion category: there is another non-invertible line which commutes with the Virasoro algebra, but not the extended chiral algebra [11, 12]. The TDLs in this category and their eigenvalues are given in table 2.
$(l,m)$ $(3,3)$ $(3,1)$ $(3,-1)$ $(2,0)$ $(1,1)$ $(2,2)$
$\eta$ $1$ $-\alpha$ $\alpha^{2}$ $1$ $-\alpha$ $\alpha^{2}$
$\eta^{2}$ $1$ $\alpha^{2}$ $-\alpha$ $1$ $\alpha^{2}$ $-\alpha$
$W$ $\varphi$ $\varphi$ $\varphi$ $-\varphi^{-1}$ $-\varphi^{-1}$ $-\varphi^{-1}$
$\eta W$ $\varphi$ $-\alpha\varphi$ $\alpha^{2}\varphi$ $-\varphi^{-1}$ $-\alpha\varphi^{-1}$ $-\alpha^{2}\varphi^{-1}$
$\eta^{2}W$ $\varphi$ $\alpha^{2}\varphi$ $-\alpha\varphi$ $-\varphi^{-1}$ $-\alpha^{2}\varphi^{-1}$ $-\alpha\varphi^{-1}$
$N$ $\sqrt{3}$ 00 00 $-\sqrt{3}$ 00 00
$WN$ $\varphi\sqrt{3}$ 00 00 $\varphi^{-1}\sqrt{3}$ 00 00
Table 2: The fusion subcategory of the monster CFT inherited from the three-state Potts model. Here $\alpha=(-1)^{1/3}$, and $\varphi=\frac{1+\sqrt{5}}{2}$.
We note that the only characters appearing in the monstralizer decomposition of $J$ in this case correspond to the chiral primaries of the three-state Potts model (rather than the Virasoro primaries), and so the TDL eigenvalues are only given for these. The lines labeled $W$, $\eta W$, $\eta^{2}W$, $N$, and $WN$ are non-invertible. Once again, using the Hecke operator for vector-valued modular forms and equation 2.6, we are able to compute $q$-expansions for the defect McKay-Thompson series associated to each line. $Z^{\mathcal{\eta}}$ and $Z^{\mathcal{\eta}^{2}}$ are the usual McKay-Thompson series for the $\mathbb{Z}_{3A}$ conjugacy class; the other series are novel.
2.3 $\quantity(\mathbb{Z}_{4\text{A}},4.2^{22}.Co_{3})$: $\mathbb{Z}_{4}$ parafermion
Our next pair of monstralizer VOAs are the $\mathbb{Z}_{4}$ parafermion CFT and its commutant in $V^{\natural}$. In this case, the Verlinde lines do exhaust the fusion category, and are given along with their eigenvalues in table 3 [11].
$(l,m)$ $(4,4)$ $(4,2)$ $(4,0)$ $(4,-2)$ $(3,3)$ $(3,1)$ $(3,-1)$ $(1,1)$ $(2,0)$ $(2,2)$
$\theta$ $1$ $-1$ $1$ $-1$ $i$ $-i$ $i$ $-i$ $1$ $-1$
$\theta^{2}$ $1$ $1$ $1$ $1$ $-1$ $-1$ $-1$ $-1$ $1$ $1$
$\theta^{3}$ $1$ $-1$ $1$ $-1$ $-i$ $i$ $-i$ $i$ $1$ $-1$
$X$ $\sqrt{3}$ $\sqrt{3}i$ $-\sqrt{3}$ $-\sqrt{3}i$ $\beta$ $\beta^{3}$ $-\beta$ $-\beta^{3}$ 00 00
$\theta X$ $\sqrt{3}$ $-\sqrt{3}i$ $-\sqrt{3}$ $\sqrt{3}i$ $\beta^{3}$ $\beta$ $-\beta^{3}$ $-\beta$ 00 00
$\theta^{2}X$ $\sqrt{3}$ $\sqrt{3}i$ $-\sqrt{3}$ $-\sqrt{3}i$ $-\beta$ $-\beta^{3}$ $\beta$ $\beta^{3}$ 00 00
$\theta^{3}X$ $\sqrt{3}$ $-\sqrt{3}i$ $-\sqrt{3}$ $\sqrt{3}i$ $-\beta^{3}$ $-\beta$ $\beta^{3}$ $\beta$ 00 00
$Y$ $2$ $2$ $2$ $2$ 00 00 00 00 $-1$ $-1$
$\theta Y$ $2$ $-2$ $2$ $-2$ 00 00 00 00 $-1$ $1$
Table 3: The fusion subcategory of the monster CFT inherited from the $\mathbb{Z}_{4}$ parafermion CFT. Here $\beta=(-1)^{1/4}$.
As in the previous cases, the primaries appearing in the decomposition at the level of characters are the primaries of the extended chiral algebra (i.e. not all Virasoro primaries). Once again there is a Hecke operator relating the parafermion characters to those of the commutant, in this case $\mathsf{T}_{23}$. All lines besides powers of $\theta$ are non-invertible and so give novel defect McKay-Thompson series; $\theta$ implements the global symmetry corresponding to the $\mathbb{Z}_{4A}$ conjugacy class of the monster.
2.4 $\quantity(D_{3\text{A}},Fi_{23})$: $\mathcal{W}_{D_{3\text{A}}}$ algebra
Finally, we consider the monstralizer pair $\quantity(D_{3\text{A}},Fi_{23})$. In the uplift to VOAs, $D_{3A}$ becomes $\mathcal{W}_{D_{3\text{A}}}$, the chiral algebra of which has been studied in [13, 14, 15]. Comparatively little is known about this VOA: in particular, we are not aware of any attempt to derive the full fusion category of TDLs in the theory. Fortunately, the characters of the theory may be decomposed into minimal model characters and consequently the modular $S$ and $T$ matrices may be derived [6]. We may therefore obtain the Verlinde lines of the theory using Verlinde’s formula [16], but there may be additional TDLs not found by this method. The TDLs found in this way and their eigenvalues are given in table 4.
$h$ 00 $\frac{1}{7}$ $\frac{5}{7}$ $\frac{2}{5}$ $\frac{19}{35}$ $\frac{4}{35}$
$A$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$
$B$ $\gamma^{-1}$ $\alpha^{-1}$ $\beta^{-1}$ $\gamma^{-1}$ $\alpha^{-1}$ $\beta^{-1}$
$W$ $\varphi$ $\varphi$ $\varphi$ $-\varphi^{-1}$ $-\varphi^{-1}$ $-\varphi^{-1}$
$AW$ $\alpha\varphi$ $\beta\varphi$ $\gamma\varphi$ $-\alpha\varphi^{-1}$ $-\beta\varphi^{-1}$ $-\gamma\varphi^{-1}$
$BW$ $\gamma^{-1}\varphi$ $\alpha^{-1}\varphi$ $\beta^{-1}\varphi$ $-\gamma^{-1}\varphi^{-1}$ $-\alpha^{-1}\varphi^{-1}$ $-\beta^{-1}\varphi^{-1}$
Table 4: The fusion subcategory of the monster CFT inherited from the $\mathcal{W}_{D_{3\text{A}}}$ CFT. Here $\alpha=\frac{\cos(3\pi/14)}{\sin(\pi/7)}$, $\beta=-2\sin(\frac{3\pi}{14})$, $\gamma=2\sin(\frac{\pi}{14})$, and $\varphi=\frac{1+\sqrt{5}}{2}$, and the primaries are labeled by their (left-moving) conformal dimensions.
All the Verlinde lines are non-invertible. This is the first case we have considered where the characters of the VOAs forming the monstralizer pair are not known to be related by a Hecke operator; however, the characters of the commutant theory solve a sixth order modular linear differential equation [6], and one may pair them with the correct $\mathcal{W}_{D_{3\text{A}}}$ characters by requiring that
$J(\tau)=\sum_{i}\chi^{(W_{D_{3\text{A}}})}_{i}(\tau)\chi^{(Fi_{23})}_{i}(\tau).$ (2.8)
This allows us to compute $q$-expansions for the defect McKay-Thompson series.
3 Invariance groups
Now that we have collected a number of defect McKay-Thompson series, we wish to find their modular invariance groups. We do so in two parts: first describing a general method for determining the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and then considering whether the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{R})$ might be larger.
3.1 In $\operatorname{\textsl{SL}}(2,\mathbb{Z})$
Turning to the invariance groups in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, we begin by exposititing the method used by Lin and Shao to compute the invariance group of the defect McKay-Thompson series found in [5]. As we shall see, this method is not suited to the more general case, and so we develop an algorithm for determining the modular invariance group of any defect McKay-Thompson series in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$.
3.1.1 Modular $S$ and $T$ matrices
In [5], Lin and Shao showed that the defect McKay-Thompson series $Z^{\mathcal{D}}(\tau)$ (corresponding to the $\mathcal{D}$ line in the Ising category) is invariant under the action of the congruence subgroup of $\operatorname{\textsl{PSL}}(2,\mathbb{Z})$ called $16D^{0}$ in the notation of [17]. To see this, consider the larger congruence subgroup $\Gamma_{0}(8)$, where
$\Gamma_{0}(N)=\quantity{\matrixquantity(a&b\\ c&d)\in\operatorname{\textsl{PSL}}(2,\mathbb{Z})\ \bigg{|}\ c\equiv 0\mod N}.$ (3.1)
Lin and Shao pointed out that $\Gamma_{0}(8)$ is generated by three matrices (say $T$, $G_{1}$, $G_{2}$) satisfying
$\begin{gathered}Z^{\mathcal{D}}(G_{1}\cdot\tau)=-Z^{\mathcal{D}}(\tau),\mbox{\quad\quad}Z^{\mathcal{D}}(G_{2}\cdot\tau)=Z^{\mathcal{D}}(\tau)\\ \mbox{\quad and\quad}Z^{\mathcal{D}}(T\cdot\tau)=Z^{\mathcal{D}}(\tau+1)=Z^{\mathcal{D}}(\tau).\end{gathered}$ (3.2)
Thus, $Z^{\mathcal{D}}(\tau)$ is invariant under any element of $\Gamma_{0}(8)$ which may be written as a word in $T$, $G_{1}$, and $G_{2}$ containing an even number of $G_{1}$. They went on to show that this subgroup is precisely $16D^{0}$, a level 16, genus zero subgroup of $\operatorname{\textsl{PSL}}(2,\mathbb{Z})$ [17].
One might hope that Lin and Shao’s procedure could be adapted to find the invariance groups in $SL(2,\mathbb{Z})$ of other defect McKay-Thompson series; to that end, we note that 8 is the lowest value of $n$ for which the modular tranformation $ST^{n}S$ fixes $Z^{\mathcal{D}}(\tau)$ up to a sign. This suggests a general procedure for finding the invariance groups: starting with a defect McKay-Thompson series $Z^{\mathcal{L}}(\tau)$, find the lowest value of $n$ such that
$Z^{\mathcal{L}}(ST^{n}S\cdot\tau)=\pm Z^{\mathcal{L}}(\tau).$ (3.3)
Then take a set of generators for the group $\Gamma_{0}(n)$ (say $\quantity{G_{i}}$) and compute $Z^{\mathcal{L}}(G_{i}\cdot\tau)$ for each. Hopefully, one obtains relations like (3.2), using which one can describe the invariance group in $\Gamma_{0}(n)$. Several difficulties prevent this. For one, this procedure does not guarantee that there are no other elements of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ leaving $Z^{\mathcal{D}}(\tau)$ invariant (in other words, the invariance group of $Z^{\mathcal{L}}$ in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ might be larger than its invariance group in $\Gamma_{0}(n)$); moreover, for more general defect McKay-Thompson series, the analogous relations to (3.2) are complicated and do not lend themselves to a simple description of the invariance group in terms of restrictions on words in the generators. In the next section we describe a completely general method for computing the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ of a defect McKay-Thompson series.
3.1.2 Coset permutations
Our general method for finding the invariance groups in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ begins with the following observation: from equations 2.6 and 2.3, the defect McKay-Thompson series $Z^{\mathcal{L}}(\tau)$ is invariant under the modular transformation $M$ if
$\rho(M)^{\dagger}D_{\mathcal{L}}\rho(M)=D_{\mathcal{L}},$ (3.4)
where $\rho$ is the representation of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ generated by the modular $S$ and $T$ matrices of $\mathcal{A}$. Moreover, this condition is necessary provided that the characters $\chi_{i}^{(\mathcal{A})}(\tau)$ appearing in the monstralizer decomposition (2.2) are linearly independant. In all the cases under consideration in this work, the various characters either (i) have different leading powers of $q$ or (ii) are equal. In case (i), the characters must obviously be linearly independant. In case (ii) the Hecke images of the redundant characters will also be equal, which means we may represent the McKay-Thompson in an alternate way. Suppose
$\chi_{1}^{(\mathcal{A})}(\tau)=\chi_{2}^{(\mathcal{A})}(\tau).$ (3.5)
then $\chi_{1}^{(\mathcal{B})}(\tau)$ and $\chi_{2}^{(\mathcal{B})}(\tau)$ are equal as well. Therefore, write
$\vec{\tilde{\chi}}^{(\mathcal{A})}(\tau)=\matrixquantity(\sqrt{2}\chi_{1}^{(\mathcal{A})}(\tau)\\ \chi_{3}^{(\mathcal{A})}(\tau)\\ \chi_{4}^{(\mathcal{A})}(\tau)\\ ...)\mbox{\quad and\quad}\vec{\tilde{\chi}}^{(\mathcal{B})}(\tau)=\matrixquantity(\sqrt{2}\chi_{1}^{(\mathcal{B})}(\tau)\\ \chi_{3}^{(\mathcal{B})}(\tau)\\ \chi_{4}^{(\mathcal{B})}(\tau)\\ ...).$ (3.6)
This furnishes our alternate representation of $Z^{\mathcal{L}}(\tau)$:
$Z^{\mathcal{L}}(\tau)=\vec{\chi}^{(\mathcal{A})}(\tau)^{\intercal}D_{\mathcal{L}}\vec{\chi}^{(\mathcal{B})}(\tau)=\vec{\tilde{\chi}}^{(\mathcal{A})}(\tau)^{\intercal}\tilde{D}_{\mathcal{L}}\vec{\tilde{\chi}}^{(\mathcal{B})}(\tau)$ (3.7)
where $\tilde{D}_{\mathcal{L}}=\operatorname{\text{diag}}\quantity(\frac{1}{2}(L_{1}+L_{2}),L_{3},L_{4},...)$. We pause to emphasize that equation 3.7 should be understood as a statement about modular functions, not about CFTs. The right hand side of (3.7) should not be interpreted as a “true” defect McKay-Thompson series: we have not constructed a new CFT with characters $\vec{\tilde{\chi}}^{(\mathcal{A})}$. Rather, (3.7) is merely a convenient way of rewriting $\mathcal{Z}^{\mathcal{L}}(\tau)$ in terms of linearly independent objects. Nonetheless, one may check that $\vec{\tilde{\chi}}^{(\mathcal{A})}(\tau)$ and $\vec{\tilde{\chi}}^{(\mathcal{B})}(\tau)$ in fact transform linearly under modular transformations, with “modified” $S$ and $T$-matrices
$\begin{gathered}\tilde{T}^{(\mathcal{A})}=\operatorname{\text{diag}}(T^{(\mathcal{A})}_{11},T^{(\mathcal{A})}_{33},T^{(\mathcal{A})}_{44},...)\\ \tilde{S}^{(\mathcal{A})}_{ij}=\begin{cases}\frac{1}{4}\quantity(S^{(\mathcal{A})}_{11}+S^{(\mathcal{A})}_{12}+S^{(\mathcal{A})}_{21}+S^{(\mathcal{A})}_{22})&i=j=1\\ \frac{1}{2}\quantity(S^{(\mathcal{A})}_{1j}+S^{(\mathcal{A})}_{2j})&i=1,\ j>1\\ \frac{1}{2}\quantity(S^{(\mathcal{A})}_{i1}+S^{(\mathcal{A})}_{i2})&i>1,\ j=1\\ S^{(\mathcal{A})}_{(i+1)(j+1)}&i,j>1\end{cases}\\ \tilde{T}^{(\mathcal{B})}=T^{(\mathcal{A})*}\mbox{\quad and\quad}\tilde{S}^{(\mathcal{B})}=S^{(\mathcal{A})*}.\end{gathered}$ (3.8)
Thus, the problem of finding the modular invariance group of $D_{\mathcal{L}}$ transforming in the representation generated by $S^{(\mathcal{A})}$ and $T^{(\mathcal{A})}$ is identical to that of finding the modular invariance group of $\tilde{D}_{\mathcal{L}}$ transforming in the representation generated by $\tilde{S}^{(\mathcal{A})}$ and $\tilde{T}^{(\mathcal{A})}$. This prodecure allows us to remove the redundant characters, leaving only a linearly independent list behind. Going forward we will have done so, and it may be assumed that the different elements of $\vec{\chi}^{(\mathcal{A})}(\tau)$ and $\vec{\chi}^{(\mathcal{B})}(\tau)$ are linearly independant, in particular having different leading powers of $q$ (though again we note that these elements are no longer necessarily characters of any RCFT, and should be thought of simply as modular functions which we have chosen for convenience).
Next, let $\Gamma$ be the invariance group of $Z^{\mathcal{L}}(\tau)$ in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$. We make the following claim: $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ conjugates of $D_{\mathcal{L}}$ (i.e. matrices of the form $\rho(M)^{\dagger}D_{\mathcal{L}}\rho(M)$, $M\in\operatorname{\textsl{SL}}(2,\mathbb{Z})$) are in one-to-one correspondence with the right cosets of $\Gamma$. To see this, let $G_{1}$ and $G_{2}$ be two elements of $SL(2,\mathbb{Z})$. Note that
$\rho(G_{1})^{\dagger}D_{\mathcal{L}}\rho(G_{1})=\rho(G_{2})^{\dagger}D_{\mathcal{L}}\rho(G_{2})$ (3.9)
[…]
Thus, the $G_{1,2}$-conjugates of $D$ are equal if and only if $G_{1}$ and $G_{2}$ lie in the same right coset of $\Gamma$, so distinct matrices of the form $\rho(M)^{\dagger}D_{\mathcal{L}}\rho(M)$ are indeed in one-to-one correspondence with right cosets of $\Gamma$. Moreover, the permutation of the right cosets induced by a given element of $\Gamma$ acting from the right is the same as the permutation of the $\Gamma$-conjugates of $D$ induced by conjugation by that element.
The utility of this correspondence is in the following fact: a finite-index subgroup of a finitely generated group is entirely determined by the action of the generators on its right cosets, meaning the permutation of the right cosets induced by the action of each generator from the right. This means that, if we know the possible $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ conjugates of $D_{\mathcal{L}}$, we can find the invariance group by conjugating each one by $\rho(S)$ and $\rho(T)$, and recording how each generator permutes the conjugates. To see this, begin with $D_{\mathcal{L}}$ and successively conjugate with $\rho(S)$ and $\rho(T)$, building a directed graph with two edges exiting each node, until the graph closes on itself. This is the Cayley graph corresponding to the action of $SL(2,\mathbb{Z})$ on the cosets; the graph will contain as many nodes as there are distinct $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ conjugates of $D_{\mathcal{L}}$ (which is the same as the number of cosets of $\Gamma$, a.k.a. the index of $\Gamma$ in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$). Then, find a generating set for the fundamental group of the directed graph. Each of the elements of the generating set is a sequence of edges labelled with generators of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ and so may be translated into an element of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$; these elements then generate $\Gamma$. In practice, once we have built the Cayley graph, we may number the nodes arbitrarily and find the permutation of cosets induced by $S$ (resp. $T$) by following the chain of consecutive $S$ (resp. $T$) edges. Given the permutations, the Sage mathematics system provides a built-in method for determining the invariance group [18]. An example graph is shown in figure 1.
1 2 3 5 6 7 8 4 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 $S$$T$
[…]
Finally, we note that the permutations induced by each generator uniquely fix a subgroup of $SL(2,\mathbb{Z})$ irrespective of any relabeling of the right cosets. For a given defect $\mathcal{L}$, the twined, twisted, and twisted + twined partition functions are simply related by acting by $S$ and $T$, or equivalently by conjugating $D_{\mathcal{L}}$ by $\rho(S)$ and $\rho(T)$, which amounts to relabeling the nodes of the directed graph. Therefore the twined, twisted, and twisted + twined partition functions share the same invariance group in $SL(2,\mathbb{Z})$.
3.2 In $\operatorname{\textsl{SL}}(2,\mathbb{R})$: Atkin-Lehner involutions
Some McKay-Thompson series have invariance groups in $SL(2,\mathbb{Z})$ which are not genus zero. For these series, one must include elements of $SL(2,\mathbb{R})$ in order to obtain a genus zero invariance group [4, 19]. One might wonder if defect McKay-Thompson series are the same way: perhaps one must consider $\operatorname{\textsl{SL}}(2,\mathbb{R})$ transformations to obtain a genus zero invariance group. On the other hand, if the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$ is already genus zero, the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{R})$ must be as well, since the genus of a subgroup is always greater than or equal to the genus of the supergroup [17]. This suggests that we check, for each of our TDLs $\mathcal{L}$ whose invariace group in $SL(2,\mathbb{Z})$ is not genus zero, whether there are (non-integral) elements of $\operatorname{\textsl{SL}}(2,\mathbb{R})$ which leave $Z^{\mathcal{L}}(\tau)$ invariant.
In general, since $\operatorname{\textsl{SL}}(2,\mathbb{R})$ is not finitely generated, this is an intractable problem; that said, there is a particular class of $SL(2,\mathbb{R})$ transformations which we are well-motivated to consider: the Atkin-Lehner involutions. Conway and Norton [1] argued that the invariance groups of the McKay-Thompson series have the form $\Gamma_{0}(N)+e,f,g,...$, where $N$ is the level of the invariance group as a congruence subgroup of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and $+e,f,g...$ signifies that we may have to adjoin certain cosets of $\Gamma_{0}(N)$, labeled by integers $e,f,g$ which exactly divide $N$ ($e\mathrel{\|}N$ means $e\mid N$ and $(e,N/e)$ are relatively prime). Since we may determine the level of $\Gamma$ even if we only know the invariance group in $SL(2,\mathbb{Z})$, we may assume for the moment that the defect McKay-Thompson series behave similarly, and explicitly check whether any of these cosets fix $Z_{\mathcal{L}}(\tau)$. If not, we know that either the invariance group in $SL(2,\mathbb{R})$ does not extend that in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, or the defect McKay-Thompson series have invariance groups in $SL(2,\mathbb{R})$ of a different form from those of the “normal” McKay-Thompson series.
To perform the checks, we need formulae for the characters appearing in the monstralizer decompositions (2.2). Fortunately, in nearly all of the cases discussed in section 2, one set of characters (those of the “small” RCFT) is given in terms of minimal model characters, and the other set is given either by the action of a vector-valued Hecke operator on the first. Only in the case with monstralizer pair $\quantity(D_{3\text{A}},Fi_{23})$ are the characters related by a modular linear differential equation without a (known) solution in terms of Hecke operators. In all other cases, we may therefore express $Z_{\mathcal{L}}(\tau)$ as an infinite sum. Next, the $\operatorname{\textsl{SL}}(2,\mathbb{R})$-transformed characters are computed from the untransformed ones by Poisson resummation, allowing us to check whether a particular transformation fixes $Z_{\mathcal{L}}(\tau)$. For each series which is not genus zero, we may obtain the level using our method for finding the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and then check each of the associated Atkin-Lehner involutions to see if the series is invariant. Doing so, we do not find that any of the Atkin-Lehner involutions extend the invariance groups in any case (apart from $D_{3A}$, which we cannot check with our methods).
4 Results
In this section we present our main results: the invariance groups of all of the (defect) McKay-Thompson series described in section 2. Table 5 gives, for each of the TDLs studied,
  1. 1.
    the TDL’s label (cf. section 2),
  2. 2.
    whether the TDL is invertible (and so whether its McKay-Thompson series is of the traditional or “defect” type),
  3. 3.
    the level of the invariance group as a congruence subgroup of $\operatorname{\textsl{PSL}}(2,\mathbb{Z})$,
  4. 4.
    the genus of the invariance group,
  5. 5.
    the index of the invariance group in $SL(2,\mathbb{Z})$, and
  6. 6.
    the label of the invariance group, in the notation of [17], if available. (*) means that the group has genus greater than 3, in which case [17] does not give a label for it. For these groups, generators are given in appendix.
TDL Invertible? Level Genus Index Label
Ising
$\eta$ Yes 2 0 3 $2B^{0}$
$\mathcal{D}$ No 16 0 24 $16D^{0}$
Three-state Potts
$\eta$ Yes 3 0 4 $3B^{0}$
$\eta^{2}$ Yes 3 0 4 $3B^{0}$
$W$ No 5 0 12 $5D^{0}$
$\eta W$ No 15 1 48 $15G^{1}$
$\eta^{2}W$ No 15 1 48 $15G^{1}$
$N$ No 15 1 48 $15G^{1}$
$WN$ No 15 1 48 $15G^{1}$
$\mathbb{Z}_{4}$ parafermion
$\theta$ Yes 4 0 6 $4B^{0}$
$\theta^{2}$ Yes 2 0 3 $2B^{0}$
$\theta^{3}$ Yes 4 0 6 $4B^{0}$
$X$ No 48 7 192 *
$\theta X$ No 48 7 192 *
$\theta^{2}X$ No 48 7 192 *
$\theta^{3}X$ No 48 7 192 *
$Y$ No 6 0 12 $6F^{0}$
$\theta Y$ No 12 0 24 $12E^{0}$
$\mathcal{W}_{D_{3\text{A}}}$
$A$ No 7 0 24 $7E^{0}$
$B$ No 7 0 24 $7E^{0}$
$W$ No 5 0 12 $5D^{0}$
$AW$ No 35 13 288 *
$BW$ No 35 13 288 *
Table 5: Our main results: the invariance groups associated to the TDLs in the Ising, three-state Potts, $\mathbb{Z}_{4}$ parafermion, and $\mathcal{W}_{D_{3\text{A}}}$ subcategories of the monster CFT.
The invariance groups for the invertible lines are genus zero as expected, but many of the invariance groups for the defect McKay-Thompson series are seen to have nonzero genus.
For reference, we also provide $q$ expansions for selected defect McKay-Thompson series.  [^1] These $q$ expansions were constructed using the relationships between the characters of the communtant pairs described in section 2. Many of these relationships are cataloged in [6] – however, some of the $q$-expansions for the commutant characters in that paper contain typos; to recreate these results the vector-valued Hecke operators and MLDEs should be used to recover the commutant characters as appropriate. For the Ising subcategory, the defect McKay-Thompson series associated to the Kramers-Wannier defect $\mathcal{D}$ is
$Z^{\mathcal{D}}(q)=\sqrt{2}\quantity(\frac{1}{q}+91886q+8498776q^{2}+301112552q^{3}+...)$ (4.1)
which agrees with the $q$-expansion presented in [5]. All other lines in the subcategory are invertible and so correspond to “typical” McKay-Thompson series.
For the three-state Potts subcategory, the defect McKay-Thompson series associated to the line $W$ is
$Z^{W}(q)={}$ $\frac{1}{2}\quantity(\frac{1}{q}+196884q+21493760q^{2}+864299970q^{3}+...)$ (4.2)
 $\mbox{\quad\quad}+\frac{\sqrt{5}}{2}\quantity(\frac{1}{q}-78794q-9307840q^{2}-381508370q^{3}+...).$
For the $\mathbb{Z}_{4}$ parafermion subcategory, the defect McKay-Thompson series associated to the line $Y$ is
$Z^{Y}(q)=\frac{2}{q}+181547q+19336938q^{2}+....$ (4.3)
For the $\mathcal{W}_{D_{3A}}$ subcategory, the defect McKay-Thompson series associated to the line $A$ is
$Z^{A}(q)={}$ $\cos(\frac{3\pi}{14})\csc(\frac{\pi}{7})\quantity(\frac{1}{q}-9618984q^{2}+...)$ (4.4)
 $\mbox{\quad\quad}+(36570q+5104206q^{2}+229935232q^{3}+...)$
 $\mbox{\quad\quad}+\frac{1}{2}\csc(\frac{\pi}{14})\quantity(176130q+7900063q^{2}+663713370q^{3}+...)$
 $\mbox{\quad\quad}+\cos(\frac{\pi}{7})\csc(\frac{\pi}{14})\quantity(-88362q-384310822q^{3}+...).$
5 Conclusions
Using the monstralizer pairs of CFTs cataloged in [6], we have found a number of topological defect lines (TDLs) appearing in the monster CFT. For each of these we have computed the modular invariance group of the associated defect McKay-Thompson series in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and given (partial) evidence that the invariance group in $SL(2,\mathbb{R})$ is not larger. Thus, we conclude that the invariance groups of defect McKay-Thompson series either
  • do not generically satisfy the genus-zero property, or
  • differ fundamentally from the invariance groups of “normal” McKay-Thompson series in that they contain non-integral elements of $SL(2,\mathbb{R})$ other than Atkin-Lehner involutions.
In the future, we may wish to understand whether the noninvertible lines which are genus zero fit some pattern, as well as how these results fit into the physical interpretation of the genus zero property proposed in [19, 20].
6 Acknowledgements
This work was supported by the National Science Foundation Grant PHY-2310635. JH thanks S.-H. Shao and G. Moore for helpful conversations.
Appendix A Generators for the invariance groups of high genus
Here we present generators for the invariance groups in table 5 of genus greater than 3, since the information provided in the table is not sufficient to uniquely determine them. The invariance groups of the TDLs of the $\mathbb{Z}_{4}$ parafermion subcategory labeled $X$, $\theta X$, $\theta^{2}X$, and $\theta^{3}X$ are identical, and are generated by
[…]
$\smallmatrixquantity(53&-6\\ 168&-19),\ \smallmatrixquantity(67\hfil&-9\\ 216&-29),\ \smallmatrixquantity(163&-24\\ 360&-53),\smallmatrixquantity(667&-101\\ 984&-149),\ \smallmatrixquantity(25&-4\\ 144&-23),\smallmatrixquantity(265&-49\\ 384&-71),\ \smallmatrixquantity(263&-50\\ 384&-73),$
$\smallmatrixquantity(19&-4\\ 24&-5),\ \smallmatrixquantity(53\hfil&-14\\ 72&-19),\ \smallmatrixquantity(265&-73\\ 432&-119),\ \smallmatrixquantity(139&-39\\ 360&-101),\ \smallmatrixquantity(169&-49\\ 576&-167),\ \smallmatrixquantity(169&-50\\ 240&-71),\ \smallmatrixquantity(53&-19\\ 120&-43),$
$\smallmatrixquantity(73&-27\\ 192&-71),\ \smallmatrixquantity(307\hfil&-120\\ 504&-197),\ \smallmatrixquantity(913&-358\\ 1344&-527),\ \smallmatrixquantity(409&-161\\ 912&-359),\ \smallmatrixquantity(361&-147\\ 528&-215),\ \smallmatrixquantity(317&-130\\ 456&-187),$
$\smallmatrixquantity(121&-50\\ 288&-119),\ \smallmatrixquantity(265\hfil&-121\\ 576&-263),\ \smallmatrixquantity(25&-12\\ 48&-23),\ \smallmatrixquantity(211&-117\\ 312&-173),\ \smallmatrixquantity(77&-43\\ 120&-67),\ \smallmatrixquantity(121&-75\\ 192&-119),\ \smallmatrixquantity(-1&0\\ 0&-1)\big{\}}.$
The invariance groups associated to the TDLs $AW$ and $BW$ in the $\mathcal{W}_{D_{3\text{A}}}$ subcategory are identical, and are generated by
[…]
$\smallmatrixquantity(281&-13\\ 735&-34),\ \smallmatrixquantity(41\hfil&-2\\ 595&-29),\ \smallmatrixquantity(176&-9\\ 665&-34),\ \smallmatrixquantity(314&-17\\ 665&-36),\ \smallmatrixquantity(71&-4\\ 1225&-69),\ \smallmatrixquantity(314&-19\\ 595&-36),\ \smallmatrixquantity(251&-16\\ 455&-29),$
$\smallmatrixquantity(146&-11\\ 385&-29),\ \smallmatrixquantity(139\hfil&-11\\ 455&-36),\ \smallmatrixquantity(106&-9\\ 1225&-104),\ \smallmatrixquantity(139&-13\\ 385&-36),\ \smallmatrixquantity(426&-41\\ 665&-64),\ \smallmatrixquantity(169&-18\\ 385&-41),\ \smallmatrixquantity(76&-9\\ 245&-29),$
$\smallmatrixquantity(169&-22\\ 315&-41),\ \smallmatrixquantity(244\hfil&-33\\ 525&-71),\ \smallmatrixquantity(29&-5\\ 35&-6),\ \smallmatrixquantity(111&-23\\ 140&-29),\ \smallmatrixquantity(216&-47\\ 455&-99),\ \smallmatrixquantity(244&-55\\ 315&-71),\ \smallmatrixquantity(281&-64\\ 1225&-279),$
$\smallmatrixquantity(64&-15\\ 175&-41),\ \smallmatrixquantity(111\hfil&-29\\ 245&-64),\ \smallmatrixquantity(211&-57\\ 385&-104),\ \smallmatrixquantity(76&-21\\ 105&-29),\ \smallmatrixquantity(99&-29\\ 140&-41),\ \smallmatrixquantity(386&-117\\ 805&-244),\ \smallmatrixquantity(386&-121\\ 1225&-384),$
$\smallmatrixquantity(71&-23\\ 105&-34),\ \smallmatrixquantity(181\hfil&-63\\ 385&-134),\ \smallmatrixquantity(316&-111\\ 595&-209),\ \smallmatrixquantity(181&-64\\ 280&-99),\ \smallmatrixquantity(64&-25\\ 105&-41),\ \smallmatrixquantity(41&-17\\ 70&-29),\ \smallmatrixquantity(99&-43\\ 175&-76),$
$\smallmatrixquantity(561&-256\\ 1225&-559),\ \smallmatrixquantity(246\hfil&-113\\ 455&-209),\ \smallmatrixquantity(274&-131\\ 525&-251),\ \smallmatrixquantity(414&-199\\ 595&-286),\ \smallmatrixquantity(356&-187\\ 455&-239),\ \smallmatrixquantity(134&-85\\ 175&-111),\ \smallmatrixquantity(-1&0\\ 0&-1)\big{\}}.$